OpenStudy (anonymous):

Peggy had four times as many quarters as nickels. She had $2.10 in all. How many nickels and how many quarters did she have? If the variable n represents the number of nickels, then which of the following expressions represents the number of quarters? The possible answers are: n/4 n 4n n + 4

OpenStudy (whpalmer4):

let Q be the number of quarters and N be the number of nickels. Write equations to show the relationships: "four times as many quarters as nickels" \[Q = 4*N \] "$2.10 in all" Value of quarters = 25Q Value of nickels = 5N (I prefer to work in cents rather than dollars, so there are no decimals to bedevil us) \[25Q + 5N = 210\] Now solve the system of equations: \[Q = 4N\]\[25Q+5N =210\] Do you know how to do that?

OpenStudy (anonymous):

I think I do...... Let me try

OpenStudy (whpalmer4):

That's what we like to hear! :-)

OpenStudy (anonymous):

Lol

OpenStudy (whpalmer4):

Instead of the all-too-common, whiny "can't you just give me the answer, I don't have much time" :-(

OpenStudy (anonymous):

The answer is 4n. Correct?

OpenStudy (whpalmer4):

Why, yes!

OpenStudy (anonymous):

Ya... I know what ya mean

OpenStudy (anonymous):

Thanks!

OpenStudy (whpalmer4):

How about we solve the whole problem? Just for giggles?

OpenStudy (anonymous):

Just wondering..... Are you a man or woman? lol just wondering not like it matters

OpenStudy (whpalmer4):

My friends call me "hey, you", does that help? :-)

OpenStudy (anonymous):

Man?

OpenStudy (whpalmer4):

(also, I have a beard)

OpenStudy (anonymous):

Oh! Then man

OpenStudy (whpalmer4):

and I'm starring in the circus freak show :-)

OpenStudy (anonymous):

lol

OpenStudy (whpalmer4):

math problems done while you watch :-)

OpenStudy (anonymous):

:)

OpenStudy (whpalmer4):

so, we have our two equations. have you learned about solving systems of equations yet?

OpenStudy (anonymous):

I'm posting another question.

OpenStudy (whpalmer4):

let's finish this one first, then I'll help you with the other one.

OpenStudy (anonymous):

Ralph is 3 times as old as Sara. In 4 years, Ralph will be only twice as old as Sara will be then. Find Ralph's age now. If x represents Sarah's age now, which of the following expressions represents Ralph's age in four years ? 3 x 6 x 2 x + 4 3 x + 4

OpenStudy (anonymous):

I'm bad at this stuff...

OpenStudy (whpalmer4):

Ah, but I'm a great teacher, so we'll fix that :-)

OpenStudy (whpalmer4):

So here are our two equations: \[Q = 4N\]\[25Q + 5N = 210\] Do you see how I came up with both of those equations?

OpenStudy (whpalmer4):

The problem said we had 4 times as many quarters as nickels, so Q = 4N should be obvious. Also, the value of our quarters plus the value of our nickels adds up to $2.10, or 210 cents. If we have 1 quarter, that is worth 25 cents. If we have 2 quarters, 50 cents. In general, if we have Q quarters, the value of the quarters is 25 cents per quarter * the number of quarters, or 25Q. Same for nickels, except there it is 5N. So our value of coins is 25Q + 5N = 210.

OpenStudy (whpalmer4):

Now, to solve this, we look at our first equation and we see that it gives Q in terms of N, specifically Q = 4N. We rewrite our second equation, replacing "Q" wherever we see it with "4N". This is the substitution step, and why this method is called (wait for it, wait for it..) substitution. \[25Q + 5N = 210\]\[25(4N) + 5N = 210\]Simplify that a bit\[100N + 5N = 210\]Simplify some more by collecting like terms\[105N = 210\]To solve that for \(N\), we just divide both sides by the number in front of \(N\): \[\frac{105N}{105} = \frac{210}{105} \]\[N = \frac{210}{105} = \]

OpenStudy (whpalmer4):

Hopefully you agree that is N = 2. Now we plug N = 2 into our formula Q = 4N Q = 4(2) = 8 So there are 8 quarters, and 2 nickels. Let's check our answer: 8*25 = 200 cents 2*5 = 10 cents ----- 210 cents (or $2.10) Our answer is correct!

OpenStudy (anonymous):

sorry i didnt reply